OpenStudy (anonymous):

1.Create your own term using a rational exponent. 2: Convert it to a radical expression. 3: Explain, in complete sentences, how the expression was converted

OpenStudy (anonymous):

So the cool thing about rational exponents is that the numerator just becomes a power and the denominator just becomes a root. Easy-peasy-parchesie. \[\Large x^\frac{a}{b} = \sqrt[b]{x^a}\]

OpenStudy (anonymous):

am i suppose to make up my own x a & b?

OpenStudy (anonymous):

Aw, yeah gurl. Rational exponent just means that it's a fraction. So make an exponent that is a fraction, and then use the rule to write it as a power and a root instead.

OpenStudy (anonymous):

oh okay. So if i was to use 61/3 then the rational exponet would be 3squareroot6?

OpenStudy (anonymous):

*6^1/3

OpenStudy (anonymous):

Yes =)

OpenStudy (anonymous):

thanks:)

OpenStudy (anonymous):

\[\Large 6^\frac{1}{3} = \sqrt[3]{6^1}\]

OpenStudy (anonymous):

:) got it!